1. define the variables used in the situation
2. translate the situation using an equality
3. represent the situation in the Cartesian plane and shade the solution set.
A parking lot has a surface of 720 m^2. Each car occupies an area of 6m^2 and each bus occupies an area of 18m^2
Please help!

Answer by rothauserc(4718)

Did your teacher give you guidance, that is, does the problem require 1 car for each bus?
:
student says maybe, so assume yes, then 6 m^2 + 18 m^2 = 24 m^2
:
consider factors of 24
:
1, 24
2, 12
3, 8
4, 6
:
consider factors of 720
:
1, 720
2, 360
3, 240
4, 180
5, 144
6, 120
8, 90
9, 80
10, 72
12, 60
15, 48
16, 45
18, 40
20, 36
24, 30
:
each set of factors of 24 represent the dimensions(length and width) of the one tile
:
each set of factors of 720 represent the dimensions(length and width) of the parking lot
:
Note 720/24 = 30 max tiles
:
Next divide the length of the parking lot by the length of the tile to work out how many tiles fit across the parking lot. Also divide the width of the parking lot by the width of the tile to work out how many tiles fit along the width of the parking lot. Once you have these two answers multiply them together.
:
So we look for dimensions in each tile pair that divide the dimensions of the parking lot since we do not want to have a fraction of a tile
:
Do all 30 tiles have to be used?
:
For example, consider the pairs 4,6 and 24,30. If we divide lengths and widths, we get 6 * 5 = 30
: